# Graphite: A GPU-Accelerated Mixed-Precision Graph Optimization Framework

### v1

@@ Abstract
We present Graphite, a GPU-accelerated non-linear graph optimization framework. It provides a CUDA C++ interface to enable the sharing of code between a real-time application, such as a SLAM system, and its optimization tasks. The framework supports techniques to reduce memory usage, including in-place optimization, support for multiple floating point types and mixed-precision modes, and dynamically computed Jacobians. We evaluate Graphite on well-known bundle adjustment problems and find that it achieves similar performance to MegBA, a solver specialized for bundle adjustment, while maintaining generality and using less memory. We also apply Graphite to global visual-inertial bundle adjustment on maps generated from stereo-inertial SLAM datasets, and observe speed ups of up to $59\times$ compared to a CPU baseline. Our results indicate that our solver enables faster large-scale optimization on both desktop and resource-constrained devices.
@@ I Introduction
To date, there have been several efforts to speed up existing optimization libraries by identifying and offloading expensive steps to GPUs. More recently, libraries such as Ceres Solver [4] have implemented GPU acceleration for the linear solver step of trust region optimization algorithms. Other work based on g2o [3] focuses on efficiently reducing the size of the linear system to be solved [7] and accelerating numerical differentiation for computing Jacobian matrices [8]. Yet, while it is possible to improve the performance of existing libraries by offloading individual steps to the GPU, it introduces additional overhead from allocating GPU memory and transferring data between steps.
@@
An alternative approach is to perform the entire optimization on the GPU. Some methods use Python [9] or a domain-specific language (DSL) [10, 11] to allow users to write simple descriptions of nonlinear optimization problems which are then used to automatically generate efficient solvers using complex code transformations. Others leverage machine learning frameworks to support batched optimization [12] and differentiable optimization [13, 14] for training learning-based tasks such as feature extraction and feature matching. In addition, many solvers are specialized for specific classes of problems such as bundle adjustment [15], achieving large speedups.
@@
However, existing GPU-accelerated solvers have several limitations which make them challenging to adapt for applications such as SLAM. They often model optimizable variables as simple numeric data types (e.g. a float, double, vector of doubles, etc.) or only provide data types for a specific optimization problem (e.g. a class representing an SE(3) transformation). Meanwhile, SLAM systems may represent variables as complex classes which consist of other classes. For example, an optimizable pose class may consist of multiple SE(3) transformations and cameras [1]. Existing solvers cannot represent these classes with a vector of numbers or predetermined types because they cannot model their complex behaviours and data dependencies. Additionally, researchers develop real-time SLAM systems in a specific language of their choice, e.g., C++, making it difficult to use DSLs or Python-based optimizers, since they require optimizable data types and functionality to be faithfully reimplemented for another language. A third challenge is that GPU memory is often limited and shared between multiple tasks such as image processing and model inference. Moreover, existing libraries may demand additional GPU memory for operations on sparse matrices and linear solver workspace allocations [16]. Even worse, embedded platforms often have no dedicated GPU memory, so the CPU and GPU must compete for the same memory.
@@
In this paper, we present Graphite, a GPU-accelerated nonlinear optimization framework for estimation problems, which implements several techniques for balancing runtime performance and GPU memory usage, to enable large-scale optimization for desktop and embedded scenarios. It allows for mixed-precision solving using 64-bit, 32-bit, and 16-bit floating-point precisions, enabling faster and more memory-efficient optimization of graphs. To further reduce memory usage, the library supports dynamically computed Jacobians for matrix-free methods, as well as an automatic differentiation method with equivalent memory overhead to analytic differentiation. In addition, the iterative linear solver is aware of the structure of the graph, bypassing explicit sparse matrix formats and their associated memory costs. To support real-time SLAM and odometry, the framework provides a CUDA C++ interface, and uses a batching model which supports in-place optimization, meaning that users can use their own data types, without having to transform them into a solver-compatible format. This avoids memory usage and runtime overhead from unnecessary copying, as well as explicit data transfers on platforms which share CPU and GPU memory.
@@
To demonstrate the effectiveness and flexibility of our framework, we evaluate our solver on bundle adjustment problems, and find it performs comparably to MegBA [15] while using up to 78% less GPU memory. We also reimplement visual-inertial bundle adjustment on the GPU inside ORB-SLAM3, which can take hundreds of seconds on the CPU for sufficiently large problems, and demonstrate a speed up of up to $59\times$ on maps generated by processing stereo-inertial SLAM datasets [17], while using under 1 GiB of GPU memory on the desktop. The code is available at https://github.com/sfu-rsl/graphite.
@@
  • We evaluate our solver across different precisions and differentiation modes, using well-known bundle adjustment and SLAM datasets.
@@ II Background
Optimizing Over Constraints: Some estimation problems can be modelled as an optimization over constraints. Each constraint represents a relationship between estimated parameters, an observation, and the error. The goal is then to find the parameters which minimize the total error across all constraints ($\mathcal{C}$). This may be formulated as
@@
Levenberg-Marquardt for Nonlinear Least Squares: Given the non-linear optimization problem in Eq. 1, the Levenberg-Marquardt algorithm (LM) can be used to iteratively find new parameters to drive the error towards a local minimum. LM combines aspects of Gauss-Newton and gradient descent to achieve better convergence with the use of a damping factor. It exploits the first-order Taylor expansion
@@
$\footnotesize(J^{\top}\Sigma^{-1}J+\lambda I)\Delta x=-J^{\top}\Sigma^{-1}r.$ (5)
This can be more simply rewritten in terms of the regularized Hessian approximation $H$ and gradient $b$ as
$H\Delta x=-b.$ (6)
Solving this linear system in each iteration yields $\Delta x$, which represents a candidate parameter step to reduce the total error. Here, $H$ is a sparse positive semi-definite block matrix.
Visual-Inertial Optimization: Given a set of keyframes $\kappa$ and $l$ map points, the goal of visual bundle adjustment is to find new parameters which better satisfy camera reprojection error constraints. The cost may be formulated as Eq. 7  [1]
$\footnotesize\sum_{j=0}^{l-1}\sum_{i\in\kappa^{j}}\rho_{\text{Huber}}(\|r_{ij}\|_{\Sigma_{ij}^{-1}}^{2})$ (7)
where $r_{ij}$ is the residual vector of the reprojection constraint between keyframe $i$ and observed map point $j$. In other words, $r_{ij}$ is the difference, in 2D coordinates, between the measured observation of a map point in a keyframe image and its modelled observation computed from the camera’s projection function using current keyframe and map point parameters. To reduce sensitivity to feature mismatches between keyframes and potential outliers, a robust Huber kernel $\rho_{\text{Huber}}$ is used [1]. Visual-inertial bundle adjustment introduces additional terms to the optimization for inertial parameters. For k+1 keyframes, this is represented as [1]
$\footnotesize\sum_{i=1}^{k}\|r_{I_{i-1,i}}\|_{\Sigma_{I_{i,i+1}}^{-1}}^{2}+\sum_{i=1}^{k}\|r_{b_{i-1,i}}\|_{\Sigma_{b_{i,i+1}}^{-1}}^{2}$ (8)
where $r_{I_{i-1,i}}$ is the inertial residual between consecutive keyframes, representing the discrepancy between preintegrated IMU measurements and the change computed from keyframe parameters. These parameters include rotation, velocity, and position. Similarly, $r_{b_{i-1,i}}$ is the accelerometer and gyroscope bias change between consecutive keyframes.
GPU Programming: Modern GPUs are based on massively parallel Single-Instruction, Multiple-Thread (SIMT) architecture, where threads execute common instructions in groups. CUDA [18] is a programming model which enables writing specialized functions, called kernels, which can be dispatched onto a GPU in order to processes a batch of similar tasks in parallel. CUDA uses a hierarchical execution model where threads are grouped into 3D thread blocks. Thread blocks are further organized into a grid. When a kernel is launched, threads and blocks are assigned identifiers which can be queried by a task to decide inputs, outputs, and behaviour. Threads can then communicate with each other using shared memory and built-in functions. Additionally, multiple kernels can be launched in different streams, where each stream is a sequence of commands, to allow their executions to overlap on the GPU.
@@ III Related Work
Although there are several versatile libraries for least squares optimization in robotics and computer vision [2, 3, 4, 19, 20, 21], they are primarily CPU-based and they generally limit GPU acceleration to the linear solver stage [4]. Our framework differs by performing the entire optimization on the GPU, to enable greater parallelization for variable and constraint operations, and avoid data transfers between intermediate steps to improve performance.
There are also a number of GPU-accelerated solvers based on domain-specific languages (DSLs) [9, 10, 11] or Python-based machine learning frameworks [12, 13, 14]. However, these approaches may result in code duplication, lack the expressiveness needed to model complex types, or involve explicit language interoperation, depending on the application. Our framework avoids this by allowing problems to be described directly using user-defined types, preventing code duplication, and enabling easier prototyping.
Several prior works also explore specialized implementations for bundle adjustment on the GPU, applying direct or iterative solving techniques [22, 23, 24, 25, 15, 26, 27]. In practice, these libraries are designed and implemented for a specific type of visual bundle adjustment which only uses camera poses, 3D points, and binary constraints. Meanwhile, SLAM and odometry systems often add several new constraints by incorporating inertial information into the optimization [1, 28], where some constraints have as many as six variables each. At the same time, they may perform other optimization tasks such as motion-only bundle adjustment and pose graph optimization [1, 28]. Our framework resolves this limitation by supporting user-defined types along with unary, binary, and n-ary GPU constraints.
@@ IV Design
Our optimization framework, based on the Levenberg-Marquardt algorithm (Section II), allows for an optimizable graph to be constructed from user-defined types and unary, binary, and n-ary constraints, to support a wide variety of optimization problems. Its design is guided by the need to minimize the amount of time needed for optimization while also using as little GPU memory as possible, which is crucial for resource-constrained systems. Next, we describe aspects of our framework that are relevant to this approach.
IV-A Overview
@@
Graphite implements graph-based optimization, where all major steps run on the GPU. It allows optimizable variables (vertices) to be represented by classes defined by the user, which avoids the need to first transform variables into a specific format for optimization, a step which is often required by other libraries. These variables can then be iteratively refined in-place on the GPU, using a batching method described in Section IV-B. As is standard with other frameworks, Graphite supports fast analytic differentiation (Section IV-C) and convenient automatic differentiation (Section IV-D) for computing Jacobian matrices used by the optimization algorithm. To perform the optimization process quickly and in a memory-efficient manner, Graphite also supports mixed-precision solving (Section IV-E).
IV-B Batching
@@
Suppose a user wants to optimize noisy 2D points so that they lie on the circumference of a circle. A typical approach to offloading work onto a GPU is to organize similar tasks into batches, and process tasks within a batch concurrently across numerous threads in order to take advantage of the massively parallel architecture. However, existing graph-based optimizers store different types of vertices and constraints indiscriminately together in the graph. This is not ideal for batching work on the GPU, because each thread processing an item (a vertex or constraint) may execute a different code path, which is detrimental to performance [29]. Therefore, to enable a GPU-friendly approach for graph-based optimization, Graphite allows users to construct an optimizable graph from collections of vertices and constraints of the same type. This enables GPU threads to process batches of identical items, which is better for performance [29]. We represent these collections using descriptors, which we describe next.
We illustrate how the user models the point optimization problem in Graphite, using mixed-precision where the user selects single precision for the descriptors, and 16-bit brain floating point for the linear system. First, the user creates a vertex descriptor for a collection of noisy points for the optimization problem, as shown in Figure 1. They specify single precision for the descriptors (Line 1) and 16-bit brain floating point for the linear system (Line 2). In Lines 4-6, they create a descriptor, reserve memory for the desired number of vertices, and add the descriptor to the graph. Then from Lines 8-10, they add a pointer for each vertex to the descriptor, and associate it with an identifier. In creating this descriptor, the user defines a batch of vertices which is processed efficiently on the GPU across many threads.
using FP = float; // vertex, constraint precision
using SP = __nv_bfloat16; // linear system precision
[…]
point_desc.reserve(num_vertices); // Reserve memory for num_vertices
graph.add_vertex_descriptor(&point_desc); // Add to graph
[…]
point_desc.add_vertex(vertex_id, &points[vertex_id]); // Add points
Figure 1: Creating a vertex descriptor for 2D points.
Next, they create a descriptor for the constraints in Figure 2. As in the previous example, they create a descriptor, reserve memory for the desired number of constraints (one constraint for each vertex), and add the descriptor to the graph (Lines 1-3). They initialize the default loss function on Line 4, and then create the desired number of constraints (Lines 6-8). When creating each constraint, the user specifies the identifiers of the vertices involved, the observation (in this case, the radius of the circle), the precision matrix (nullptr defaults to the identity matrix), the data for non-optimizable variables (not used), and the loss function parameters. Again, by constructing this descriptor, the user enables each constraint to be processed concurrently and efficiently on the GPU.
auto factor_desc = CircleFactor<FP, SP>(&point_desc); // Create descriptor
factor_desc.reserve(num_vertices); // Reserve memory for constraints
graph.add_factor_descriptor(&factor_desc); // Add descriptor to graph
const auto lf = DefaultLoss<FP, 1>(); // loss function
[…]
factor_desc.add_factor({vertex_id}, radius, nullptr, 0, lf);
Figure 2: Creating a descriptor for constraints.
@@
At compile time, the properties of each descriptor are used to specialize each CUDA kernel for operations over a specific constraint or vertex type, so that each thread carries out the same type of computation, which improves performance. For a vertex descriptor, these properties include information such as the underlying data type (class), the size of its parameter block, a method to get its parameter block, and a method to update each vertex for a new step. Figure 3 shows how the user defines the properties for the 2D point descriptor. They specify the underyling vertex data type as a $2\times 1$ Eigen [30] matrix (Line 2, Line 7), the length of the parameter block as 2 (Line 6), the method to get the parameter block (Lines 9 to 11), and the update method (Lines 13-16).
// Point definition
template <typename T> using Point = Eigen::Matrix<T, 2, 1>;
[…]
template <typename T, typename S> struct PointTraits {
static constexpr size_t dimension = 2;
using Vertex = Point<T>;
[…]
return vector_to_array<T, dimension>(vertex);
[…]
Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
vertex += d;
[…]
template <typename T, typename S>
using PointDescriptor = VertexDescriptor<T, S, PointTraits<T, S>>;
Figure 3: Properties of a vertex descriptor for a 2D point.

### v2

@@ Abstract
We present Graphite, a GPU-accelerated nonlinear least squares graph optimization framework. It provides a CUDA C++ interface to enable the sharing of code between a real-time application, such as a SLAM system, and its optimization tasks. The framework supports techniques to reduce memory usage, including in-place optimization, support for multiple floating point types and mixed-precision modes, and dynamically computed Jacobians. We evaluate Graphite on well-known bundle adjustment problems and find that it achieves similar performance to MegBA, a solver specialized for bundle adjustment, while maintaining generality and using less memory. We also apply Graphite to global visual-inertial bundle adjustment on maps generated from stereo-inertial SLAM datasets, and observe speed-ups of up to $59\times$ compared to a CPU baseline. Our results indicate that our framework enables faster large-scale optimization on both desktop and resource-constrained devices.
@@ I Introduction
To date, there have been several efforts to speed up existing optimization libraries by identifying and offloading expensive steps to GPUs. More recently, Ceres Solver [1] implemented GPU acceleration for the linear solver step of trust region optimization algorithms. Other work based on g2[18] focuses on efficiently reducing the size of the linear system to be solved [10] and accelerating numerical differentiation for computing Jacobian matrices [17]. Yet, while it is possible to improve the performance of existing libraries by offloading individual steps to the GPU, doing so introduces additional overhead from allocating GPU memory and transferring data between steps.
@@
In contrast, new GPU-based frameworks have emerged which accelerate various steps of the optimization. Some use Python [32] or a domain-specific language (DSL) [7, 19] to allow users to write simple descriptions of nonlinear optimization problems that are used to automatically generate efficient solvers using complex code transformations. Others leverage machine learning frameworks to support batched optimization [15] and differentiable optimization [34, 23, 31] for training models for tasks such as feature extraction and feature matching. In addition, many solvers are specialized for specific classes of problems such as bundle adjustment [26], achieving large speedups.
@@
However, existing GPU-accelerated solvers have several limitations which make them challenging to adapt for applications such as SLAM. They often model optimizable variables as simple numeric data types (e.g. a float, double, vector of doubles, etc.) or only provide data types for a specific optimization problem (e.g. a class representing an SE(3) transformation). Meanwhile, SLAM systems may represent variables as complex classes which consist of other classes. For example, an optimizable pose class may consist of multiple SE(3) transformations and camera references [3]. Existing solvers cannot represent these classes with a vector of numbers or predetermined data types because they cannot model their complex behaviours and data dependencies. Additionally, researchers develop real-time SLAM systems in a specific language of their choice, e.g., C++, making it difficult to use DSLs or Python-based optimizers, since they require optimizable data types and functionality to be faithfully reimplemented for another language. A third challenge is that GPU memory is often limited and shared between multiple tasks such as image processing and model inference. Moreover, existing libraries may demand additional GPU memory for operations on sparse matrices and linear solver workspace allocations [22]. Even worse, embedded platforms often have no dedicated GPU memory, so the CPU and GPU must compete for the same memory.
@@
In this paper, we present Graphite, a GPU-accelerated nonlinear least squares graph optimization framework, which implements several techniques for balancing runtime performance and GPU memory usage, to enable large-scale optimization for desktop and embedded scenarios. It allows for mixed-precision solving using 64-bit, 32-bit, and 16-bit floating-point precisions, enabling faster and more memory-efficient optimization of graphs. To further reduce memory usage, the library supports dynamically computed Jacobians for matrix-free methods, as well as an automatic differentiation method with equivalent memory overhead to analytic differentiation. In addition, the iterative linear solver is aware of the structure of the graph, bypassing explicit sparse matrix formats and their associated memory costs. To support real-time SLAM and odometry, the framework provides a CUDA C++ interface, and uses a batching model which supports in-place optimization, allowing users to optimize GPU-accessible data without first transforming it into a solver-specific format. This avoids memory usage and runtime overhead from unnecessary copying, as well as explicit data transfers on platforms which share CPU and GPU memory.
@@
To demonstrate the effectiveness and flexibility of our framework, we evaluate it on bundle adjustment problems, and find it performs comparably to MegBA [26] while using up to 78% less GPU memory. We also reimplement visual-inertial bundle adjustment on the GPU inside ORB-SLAM3, which can take hundreds of seconds on the CPU for sufficiently large problems, and demonstrate a speed up of up to $59\times$ on maps generated by processing stereo-inertial SLAM datasets [28], while using under 1 GiB of GPU memory on a desktop machine. The code is available at https://github.com/sfu-rsl/graphite.
@@
  • We evaluate our framework across different precisions and differentiation modes, using well-known bundle adjustment and SLAM datasets.
@@ II Background
Optimizing Over Constraints: Some estimation problems can be modelled as an optimization over constraints. Each constraint represents a relationship between estimated parameters, an observation, and the error. The goal is then to find the parameters that minimize the total error across all constraints ($\mathcal{C}$). This may be formulated as
@@
Levenberg-Marquardt for Nonlinear Least Squares: Given the nonlinear optimization problem in Eq. 1, the Levenberg-Marquardt algorithm (LM) can be used to iteratively find new parameters to drive the error towards a local minimum. LM modifies Gauss-Newton using a damping factor $\lambda$ to achieve better convergence. It exploits the first-order Taylor expansion
@@
$\footnotesize(J^{\top}\Sigma^{-1}J+\lambda D^{\top}D)\Delta x=-J^{\top}\Sigma^{-1}r.$ (5)
This can be more simply rewritten in terms of the regularized Hessian approximation $H_{\lambda}$ and gradient $b$ as
$\footnotesize H_{\lambda}\Delta x=-b.$ (6)
Solving this linear system in each iteration yields $\Delta x$, which is a candidate parameter step to reduce the total error. Here, $H_{\lambda}$ is a sparse symmetric positive definite block matrix [2].
Visual-Inertial Optimization: Given a set of keyframes $\kappa$ and $l$ map points, the goal of visual bundle adjustment is to find new parameters that better satisfy camera reprojection error constraints. The cost may be formulated as Eq. 7  [3]
$\footnotesize\sum_{j=0}^{l-1}\sum_{i\in\kappa^{j}}\rho_{\text{Huber}}(\|r_{ij}\|_{\Sigma_{ij}}^{2})$ (7)
where $r_{ij}$ is the residual vector of the reprojection constraint between observed map point $j$ and keyframe $i$. In other words, $r_{ij}$ is the difference, in 2D coordinates, between the computed camera projection of a map point in a keyframe image using current parameters and its measured observation. To reduce sensitivity to feature mismatches between keyframes and potential outliers, a robust Huber kernel $\rho_{\text{Huber}}$ is used [3]. Visual-inertial bundle adjustment introduces additional terms to the optimization for inertial parameters. For k+1 keyframes, this is represented as [3]
$\footnotesize\sum_{i=1}^{k}\|r_{I_{i-1,i}}\|_{\Sigma_{I_{i-1,i}}}^{2}+\sum_{i=1}^{k}\|r_{b_{i-1,i}}\|_{\Sigma_{b_{i-1,i}}}^{2}$ (8)
where $r_{I_{i-1,i}}$ is the inertial residual between consecutive keyframes, representing the discrepancy between preintegrated IMU measurements and the change computed from keyframe parameters. These parameters include rotation, velocity, and position. Similarly, $r_{b_{i-1,i}}$ is the accelerometer and gyroscope bias residual between consecutive keyframes.
GPU Programming: Modern GPUs are based on massively parallel Single-Instruction, Multiple-Thread (SIMT) architecture, where threads execute common instructions in groups. CUDA [22] is a programming model that enables writing specialized functions, called kernels, which can be dispatched onto a GPU in order to process a batch of similar tasks in parallel. CUDA uses a hierarchical execution model where threads are grouped into 3D thread blocks. Thread blocks are further organized into a grid. When a kernel is launched, threads and blocks are assigned identifiers which can be queried by a task to decide inputs, outputs, and behaviour. Threads can then communicate with each other using shared memory and built-in functions. Additionally, multiple kernels can be launched in different streams, where each stream is a sequence of commands, to allow their executions to overlap on the GPU.
@@ III Related Work
Although there are several versatile libraries for least squares optimization in robotics and computer vision [6, 18, 1, 27, 29, 20], they are primarily CPU-based and generally limit GPU acceleration to the linear solver stage [1]. Our framework differs by performing the entire optimization on the GPU, to enable greater parallelization for variable and constraint operations and avoid data transfers between intermediate steps to improve performance.
There are also a number of GPU-accelerated solvers based on domain-specific languages (DSLs) [32, 7, 19] or Python machine learning frameworks [15, 34, 23, 31]. However, these approaches may result in code duplication, lack the expressiveness needed to model complex types, or involve explicit language interoperation, depending on the application. Our framework avoids this by allowing problems to be described directly using user-defined data types, preventing code duplication, and enabling easier prototyping.
Several prior works also explore specialized implementations for bundle adjustment on the GPU, applying direct or iterative solving techniques [26, 33, 14, 9, 4, 8, 13]. While performant, these works target a specific type of visual bundle adjustment which only uses camera poses, 3D points, and binary constraints. Meanwhile, SLAM and odometry systems often add several new constraints by incorporating inertial information [3, 25], where some constraints have as many as six variables. They may also perform other optimization tasks such as motion-only bundle adjustment and pose graph optimization [3, 25]. Our framework resolves this limitation by supporting user-defined data types along with unary, binary, and n-ary GPU constraints.
@@ IV Design
Figure 1: An example Graphite workflow. The user first defines properties for their variables and constraints, then constructs a graph by creating descriptors, and finally optimizes their variables using Levenberg-Marquardt.
@@
Graphite implements nonlinear least squares graph optimization on the GPU, based on the Levenberg-Marquardt algorithm (Section II). Its design is guided by the need to minimize the amount of time required for optimization while also using as little GPU memory as possible, which is crucial for resource-constrained systems. An optimization problem is represented by a graph constructed from a set of descriptors, which are collections of optimizable vertices (variables) and unary, binary, and n-ary constraints. Descriptors provide a batching mechanism for the GPU and allow vertices and constraints to be represented with user data types that are compatible with CUDA device code (Section IV-A). This avoids the need to first transform variables into a solver-specific format for optimization, a step which is often required by other libraries. As is standard with other frameworks, Graphite supports fast analytic differentiation (Section IV-B) and convenient automatic differentiation (Section IV-C) for computing Jacobian matrices used by the optimization algorithm. To perform the optimization process quickly and in a memory-efficient manner, Graphite also supports mixed-precision solving (Section IV-D).
We show a running example of the workflow in Figure 1 for a 2D point optimization problem, where a user adjusts 2D points stored in GPU-accessible memory to lie on the boundary of a circle. They first define static properties for their 2D points and circle equation constraints. Next, they create the corresponding descriptors and construct a graph. Lastly, the user passes the graph to the Levenberg-Marquardt algorithm along with desired options, optimizing the points in-place. We discuss this process with concrete examples next.
IV-A Descriptor Batching Model
@@
A typical approach to offloading work onto a GPU is to organize similar tasks into batches, and process tasks within a batch concurrently across numerous threads in order to take advantage of the massively parallel architecture. However, existing graph-based optimizers store different types of vertices and constraints indiscriminately together in the graph. This is not ideal for batching work on the GPU, because each thread processing an item (a vertex or constraint) may branch and execute a different code path, which is detrimental to performance [22]. Therefore, to enable a GPU-friendly approach for graph-based optimization, Graphite allows users to construct an optimizable graph from collections of vertices and constraints of the same type. This enables GPU threads to process batches of identical items, which minimizes branching. We represent these collections using descriptors, which we describe next with our point optimization example.
We demonstrate how the user models the point optimization problem from Figure 1 in Graphite, where the user selects single precision for the variables and constraints, and 16-bit brain floating point for the linear system.
First, the user defines static properties for their vertices (points) and constraints (based on the circle equation), which are necessary for creating descriptors, as shown in Figure 1. At compile time, these properties are used to specialize each CUDA kernel, so that each thread carries out the same type of computation, improving performance.
@@
For a vertex descriptor, these properties include information such as the underlying data type (class), the parameter block size, a method to get its parameter block, and a method to update each vertex for a new step. 1 shows how the user defines the properties for a 2D point descriptor. They specify the underlying vertex data type as a $2\times 1$ Eigen [11] matrix (Line 2, Line 7), the length of the parameter block as 2 (Line 6), the method to get the parameter block (Lines 9 to 13), and the update method (Lines 15-18).
1// Point definition
2template <typename T> using Point = Eigen::Matrix<T, 2, 1>;
3
4// Traits for Point
5template <typename T, typename S> struct PointTraits {
6 static constexpr size_t dimension = 2;
7 using Vertex = Point<T>;
8
9 template <typename P>
10 d_fn static void parameters(const Vertex &vertex, P *parameters) {
11 Eigen::Map<Eigen::Matrix<P, dimension, 1>> params_map(parameters);
12 params_map = vertex.template cast<P>();
13 }
14
15 d_fn static void update(Vertex &vertex, const T *delta) {
16 Eigen::Map<const Eigen::Matrix<T, dimension, 1>> d(delta);
17 vertex += d;
18 }
19};
20
21// Type alias
22template <typename T, typename S>
23using PointDescriptor = VertexDescriptor<T, S, PointTraits<T, S>>;
Listing 1: Properties of a vertex descriptor for a 2D point.